OpenStudy (anonymous):

Solve the proportion. 4/10=8/a a. 4 c. 20 b. 7 d. 16

OpenStudy (jh3power):

4/10 a =8 a=8*10/4 a=20

OpenStudy (anonymous):

thanks.